OpenStudy (anonymous):

You wish to panel your den. Two walls are 12 feet long and 8 feet high. The other two walls are 8' x 8'. How many 4' by 8' panels of plywood are needed, assuming no wastage? Number of panels needed? Please help, it would be greatly appreciated. Thank-You very much.

OpenStudy (yanasidlinskiy):

Welcome to OpenStudy!!!:) (12+8+12+8) / 4 =40/4 =10